OpenStudy (anonymous):

Select the inequality that corresponds to the given graph. a number line graph with a closed circle at 3, shading to the right and an open circle at 2, shading to the left. 5x + 7 greater than or equal to 2 and 10x less than or greater to 30 6x + 7 greater than or equal to 25 or 2x - 8 < -4 4x + 10 less than or greater to 6 or 5x - 6 greater than or equal to 14 37 less than or greater to 8x - 3 less than or greater to 61

OpenStudy (texaschic101):

ok...closed circle at 3, shading to the right...means it is going to have an equal sigh in it and it will be greater then or equal to 3. An open circle at 2, shaded to the left...means no equal sign and it will be less then 2. Lets check the inequalities... 5x + 7 >= 2 5x >= 2 - 7 5x >= -5 x >= -1 (not this one...I don't even have to try the other inequality with it) 6x + 7 >= 25 6x >= 25 - 7 6x >= 18 x >= 3 (this is right....lets check the other one that goes with it 2x - 8 < -4 2x < -4 + 8 2x < 4 x < 2 (correct) I do not need to go any further...your answer is : the 2nd one :)